# Evaluating LLMs’ Mathematical and Coding Competency through Ontology-guided Interventions

### v4

@@ 4.1 Perturbed Question Generation
Instruction: Rewrite the original mathematical context below based on the #Rewrite Requirement#.
Your output should only be #Rewritten Context#.
#Original Context#: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick.
#Original Query#: What is the total inner volume of all 3 boxes?
#Rewrite Requirement#: 1. Remove some constraints or information from the original context. 2. Make sure the rewritten question can still be solved, but the answer is simpler.
#Rewritten Context#:
@@ 4.2 Filtering and Validation
Despite automatic verification, perturbed questions still have limitations, so we conduct a final human verification to refine them. Our findings show that 36% of the filtered questions needed minor rewording, 31% contained significant inaccuracies or failed the filtering, and 33% were correct as is. Thus, the final questions in More are high-quality, understandable, logically coherent, and aligned with the intended perturbation method. Human verification is performed by five PhD computer science students, with each question revised by two annotator and verified by two others.
@@ 4.3 Obtaining Final Answers of the Perturbed Questions
4.4 Statistics of More and Core
We sampled five questions from GSM8K and HumanEval and perturbed them using GPT-4 in 44 distinct perturbation categories. Following a rigorous process of filtering and validation, we retained a total of 216 and 219 perturbed questions in More and Core, respectively. We specify the detailed statistics in Appendix B and the details of the five selected question from each dataset in Appendix F and Appendix G respectively.
@@ 5.1 Evaluation Protocol
Owing to the loosely controlled format of the LLM responses to the majority of the questions, calculating accuracy through direct string matching with the annotated answer may not always be reliable. Additionally, in the context of concept analysis, curating an exhaustive list of correct answers could be intractable. For instance, the category optimize solution ((ii)) asks to further optimize the provided solution. There could be numerous distinct valid ways to optimize the given solution. To address these challenges, manual evaluation is necessary. To empirically justify this, we prompted GPT-4 for automated answer evaluation, yielding an agreement of 88.76% with human annotation on the answers of GPT-4 to More questions.
@@ 5.2 Experimental Setup
We evaluated five prominent closed- and open-sourced LLMs on our benchmark. The closed-sourced LLMs are GPT-4, GPT-3.5, and Gemini 1.5. The remaining open-sourced LLMs include one general-purpose LLM and one LLM finetuned on task-specific datasets. The general-purpose LLM is Llama3-8B-Instruct and task-specific LLMs are MetaMath-70B-V1.0 and CodeLlama-70B-Instruct for coding and maths, respectively. MetaMath-70B-V1.0 is finetuned on a mixture of datasets from Metamath (Yu et al., 2023b) and Mistral (Jiang et al., 2023) and CodeLlama-70B-Instruct is finetuned on publicly available coding and coding-related instructions Rozière et al. (2023). Model Details are specified in Appendix C.1. We listed the prompts used for these models in Appendix J. Each question is evaluated with pass@1 metric under zero-shot setting. More details in Appendix J on the evaluation settings.
@@ 5.3 Experimental Results and Analyses
Aspect | Original | Structural | Representational | ||||||||||||||||
Target | Logic | Concept | Q. Format | A. Format | Weighted | ||||||||||||||
@@
Dimension | Gran. | Reason | Compute. | Formul | Avg. | Quest. | Sol. | Crit | Avg. | Form. | Form. | Avg. | Form. | Avg. | |||||
@@
Adjust. | Adjust. | Adjust. | Adjust. | Perf. | Under. | Under. | Think. | Perf. | Change. | Comp. | Perf. | Constraint | |||||||
Maths (More) | GPT-4 | 100 | 100 | 80 | 90.91 | 60 | 78.30 | 85 | 65 | 48 | 64.62 | 90 | 60 | 84.00 | 65 | 74.21 | |||
GPT-3.5 | 80 | 75 | 27.5 | 54.55 | 25.71 | 38.68 | 55 | 45 | 12 | 35.38 | 35 | 40 | 36.00 | 5 | 35.75 | ||||
Gemini | 80 | 90 | 50 | 81.82 | 37.14 | 56.60 | 60 | 20 | 16 | 30.77 | 55 | 20 | 48.00 | 30 | 46.15 | ||||
Llama | 60 | 50 | 12.5 | 18.18 | 5.71 | 17.92 | 35 | 60 | 4 | 30.77 | 5 | 60 | 16.00 | 5 | 26.24 | ||||
Metamath | 80 | 70 | 15 | 27.27 | 11.43 | 25.47 | 30 | 25 | 4 | 18.46 | 35 | 80 | 44.00 | 20 | 21.27 | ||||
Average | 80 | 77 | 37 | 54.55 | 27.90 | 43.39 | 53 | 43 | 16.8 | 36.00 | 44 | 52 | 45.60 | 25 | 40.72 | ||||
Coding (Core) | GPT-4 | 80 | 90 | 37.5 | 46.67 | 50 | 52.29 | 65 | 80 | 44 | 61.54 | 65 | 40 | 60.00 | 55 | 56.7 | |||
GPT-3.5 | 80 | 73.68 | 35 | 40 | 29.41 | 40.74 | 60 | 75 | 40 | 56.92 | 50 | 40 | 48.00 | 45 | 47.09 | ||||
Gemini | 80 | 80 | 32.5 | 53.33 | 23.53 | 41.28 | 65 | 75 | 44 | 60.00 | 45 | 40 | 44.00 | 35 | 47.32 | ||||
Llama | 60 | 45 | 12.5 | 33.33 | 11.76 | 21.10 | 50 | 50 | 8 | 33.85 | 25 | 40 | 28.00 | 20 | 36.61 | ||||
CodeLlama | 60 | 80 | 40 | 40 | 11.76 | 38.53 | 35 | 35 | 28 | 32.31 | 40 | 0 | 32.00 | 40 | 26.34 | ||||
Average | 72 | 73.74 | 31.5 | 42.67 | 25.29 | 38.79 | 55 | 63 | 32.8 | 48.92 | 45 | 32 | 42.40 | 39.00 | 42.81 |
Mod. | Q. | R. | C. | S. | Avg. | |
Simp. | Adj. | Adj. | Man. | |||
More | GPT-4 | 100 | 77.5 | 90.91 | 71.43 | 81.13 |
GPT-3.5 | 90 | 50 | 90.91 | 40 | 58.49 | |
Gemini | 95 | 57.50 | 63.64 | 45.71 | 61.32 | |
Core | GPT-4 | 100 | 50 | 46.67 | 55.88 | 59.43 |
GPT-3.5 | 82.35 | 42.50 | 40 | 26.47 | 43.40 | |
Gemini | 70.59 | 25 | 53.33 | 26.47 | 36.79 |
Mod. | Q. | R. | C. | S. | Avg. | |
Simp. | Adj. | Adj. | Man. | |||
Self-C | GPT-4 | 95 | 87.5 | 90.91 | 65.71 | 82.08 |
GPT-3.5 | 60 | 45 | 45.45 | 25.71 | 41.51 | |
Gemini | 75 | 45 | 81.82 | 40 | 52.83 | |
POT | GPT-4 | 95 | 90 | 81.82 | 68.57 | 83.02 |
GPT-3.5 | 75 | 57.5 | 54.55 | 25.71 | 50 | |
Gemini | 90 | 60 | 63.64 | 45.71 | 61.32 |
General Performance Analysis.
@@
The results show that perturbed questions significantly challenge all models in both math and coding contexts. GPT-4’s accuracy decreased notably, as did other LLMs, with all showing a performance decline over 30 points. Notably, closed-source models outperformed open-source ones in every tested aspect. Models like CodeLlama and Metamath, fine-tuned on specific tasks, performed better in logic alteration and representational perturbations but worse in concept analysis. This suggests fine-tuning may restrict broader reasoning capacities. In general, LLMs handled logic alteration better than concept analysis, indicating their robustness in abstract reasoning yet limitations in understanding deeper mathematical concepts. GPT-4 demonstrated resilience across various question types, outshining others especially in handling different problem-solving frameworks, although it still struggled more in math than in coding in concept analysis. We include Target-wise(Level II) performance analysis in Appendix D
@@
In Table 4, providing models with the correct answer to the original question along with the prompt significantly improves their ability to solve perturbed questions, particularly in the Computational Adjustment dimension. However, performance remains weak in Symbolic Manipulation, highlighting challenges in abstract reasoning despite access to solutions. Notably, even equipped with correct answers, some models like Gemini and GPT-3.5 still fail on simpler question variants, underscoring their low sensitivity to semantic perturbations.
@@
In Table 5, different prompting techniques greatly influence model performance in logic alteration tasks. The Program-of-Thoughts technique notably boosts reasoning capabilities in closed-source models by reducing logical errors, leading to better performance in symbolic manipulation for GPT-4. Conversely, the Self-Consistency method shows only minor improvements and even a performance decline in the Gemini model, suggesting difficulties in effective in-context learning for new unseen tasks.
@@
The Formulation Adjustment dimension presents a significant challenge to both closed-source and open-source models, largely due to the demands of abstract reasoning. Instead of reasoning an number or code as answer, this involves manipulation of abstract maths and coding concepts in the logical space behind the surface of the problem. For example, in the WhatIf category, models must hypothesize outcomes by changing certain events under consistent conditions, which requires a nuanced grasp of the problem-solving framework. The Critical Thinking dimension tests a model’s ability to scrutinize relationships between pieces of information, demanding a comprehensive analysis to identify inconsistencies without a predefined solution path. This emphasizes the necessity for models to thoroughly understand and navigate through all possible avenues to effectively resolve conflicts or discrepancies. Furthermore, the Format Change dimension poses difficulties to models like ChatGPT attempt to follow these constraints but often fail to maintain the integrity of their reasoning when adapting to new formats, highlighting a lack of flexibility in handling varied task demands.
@@ 6.1 Difficulty Change by Perturbations
Category | Human Acc() | Model Acc() | Time Consumption |
Number of Reasoning Steps | |||
95.2(-4.8) | 32.0(-44.0) | 178% |
98.4(-1.6) | 44.4(-31.6) | 39% |
Reasoning Depth | |||
97.3(-2.7) | 38.4(-37.6) | 113% |
97.7(-2.3) | 42.1(-33.9) | 62% |
![[Uncaptioned image]](extracted/5694788/images/up-sign.png)
![[Uncaptioned image]](extracted/5694788/images/equal-sign.png)
@@ 7 Conclusion
Our study evaluated the robustness of several prominent Large Language Models (LLMs) in handling mathematical and coding problems. By employing an ontology for random perturbations on questions from the GSM8K and HumanEval datasets, we crafted two specialized datasets, More and Core, containing 216 and 219 questions respectively. These datasets target a broad variations of mathematical and coding problem-solving and analytical skills, resulting in notable performance drops in LLMs upon evaluation. The introduction of More and Core provides a new framework for assessing LLMs’ abilities in mathematics and coding, while also revealing their vulnerabilities in consistent reasoning across different formats. This research highlights the complex challenges that LLMs face, stressing the importance of continued exploration into their strengths and weaknesses in logical reasoning tasks. Our dataset More 222https://huggingface.co/datasets/declare-lab/GSM8k_MORE and Core 333https://huggingface.co/datasets/declare-lab/HumanEval_CORE will be publicly available online.

### v5

@@ 4.2 Filtering and Validation
Despite automatic verification, perturbed questions still have limitations, so we conduct a final human verification to refine them. Our findings show that 36% of the filtered questions needed minor rewording, 31% contained significant inaccuracies or failed the filtering, and 33% were correct as is. Thus, the final questions in GSMore and HumanEval-Core are high-quality, understandable, logically coherent, and aligned with the intended perturbation method. Human verification is performed by five PhD computer science students, with each question revised by two annotators and verified by two others.
@@ 4.3 Obtaining Final Answers of the Perturbed Questions
4.4 Statistics of GSMore and HumanEval-Core
We sampled five questions from GSM8K and HumanEval and perturbed them using GPT-4 in 44 distinct perturbation categories. Following a rigorous process of filtering and validation, we retained a total of 216 and 219 perturbed questions in GSMore and HumanEval-Core, respectively. We specify the detailed statistics in Appendix B and the details of the five selected questions from each dataset in Appendix F and Appendix G respectively.
@@ 5.1 Evaluation Protocol
Owing to the loosely controlled format of the LLM responses to the majority of the questions, calculating accuracy through direct string matching with the annotated answer may not always be reliable. Additionally, in the context of concept analysis, curating an exhaustive list of correct answers could be intractable. For instance, the category optimize solution ((ii)G27.) asks to further optimize the provided solution. There could be numerous distinct valid ways to optimize the given solution. To address these challenges, we manually evaluated the generated outputs. For HumanEval-Core, we designed test cases where possible. The evaluation is fully automated for those instances. For the remaining instances without test cases, we manually verified the answers returned by the LLMs.
Automated Evaluation.
With the recent advancements in the reasoning abilities of large language models (LLMs), we also utilize GPT-4o as an evaluator to assess the correctness of answers against the manually annotated ground truth. GPT-4o proved to be an effective evaluator, showing a high level of agreement with human judgments across various models. We report the automated evaluation results and their agreement with manual evaluation in Section 5.3.
@@ 5.2 Experimental Setup
We evaluated five prominent closed- and open-sourced LLMs on our benchmark. The closed-sourced LLMs are o1-preview, GPT-4o, GPT-4, GPT-3.5, and Gemini-1.5-pro. The remaining open-sourced LLMs include one general-purpose LLM and one LLM finetuned on task-specific datasets. The general-purpose LLM is Llama3-8B-Instruct and task-specific LLMs are MetaMath-70B-V1.0 and CodeLlama-70B-Instruct for coding and math, respectively. MetaMath-70B-V1.0 is finetuned on a mixture of datasets from Metamath (Yu et al., 2023b) and Mistral (Jiang et al., 2023) and CodeLlama-70B-Instruct is finetuned on publicly available coding and coding-related instructions Rozière et al. (2023). Model Details are specified in Appendix C.1. We listed the prompts used for these models in Appendix J. Each question is evaluated with pass@1 metric under zero-shot setting. More details are in Appendix J on the evaluation settings.
@@ 5.3 Experimental Results and Analyses
l|c|cc|ccc>cccc>ccc>ccc|c
Original Structural Representational
Logic Alteration Concept Analysis Question Format A. Format Weighted
@@
Models Seed All Gran. Reason Compute. Formul Avg. Quest. Sol. Crit. Avg. Form. Form. Avg. Form. Avg.
@@
Samp. Samp. Adjust. Adjust. Adjust. Adjust. Perf. Under. Under. Think. Perf. Change. Comp. Perf. Constraint
MATH
o1-preview 100.0 95.2 100.0 92.5 92.3 77.1 88.9 95.0 95.0 56.0 80.0 95.0 80.0 92.0 90.0 86.7
GPT-4o 100.0 95.1 100.0 92.5 81.8 71.4 85.8 95.0 90.0 48.0 75.4 100.0 80.0 96.0 85.0 83.8
GPT-4 100.0 94.6 100.0 80.0 90.9 60.0 78.3 85.0 65.0 48.0 64.6 90.0 60.0 84.0 65.0 74.2
GPT-3.5 80.0 78.9 75.0 27.5 54.6 25.7 38.7 55.0 45.0 12.0 35.4 35.0 40.0 36.0 5.0 35.8
Gemini 80.0 91.7 90.0 50.0 81.8 37.1 56.6 60.0 20.0 16.0 30.8 55.0 20.0 48.0 30.0 46.2
Llama3 60.0 80.7 50.0 12.5 18.2 5.7 17.9 35.0 60.0 4.0 30.8 5.0 60.0 16.0 5.0 26.2
Metamath 80.0 82.3 70.0 15.0 27.3 11.4 25.5 30.0 25.0 4.0 18.5 35.0 80.0 44.0 20.0 21.3
Average 85.7 88.4 83.6 52.9 63.8 41.2 56.0 65.0 57.1 26.9 47.9 59.3 60.0 59.4 42.9 53.5
CODE
o1-preview 100.0 92.4 95.0 62.5 46.7 80.0 71.8 95.0 85.0 64.0 80.0 75.0 80.0 76.0 95.0 76.8
GPT-4o 80.0 90.2 90.0 57.5 53.3 74.3 68.1 80.0 95.0 56.0 75.4 65.0 60.0 64.0 70.0 70.0
GPT-4 80.0 76.5 90.0 37.5 46.7 50.0 52.3 65.0 80.0 44.0 61.5 65.0 40.0 60.0 55.0 56.7
GPT-3.5 80.0 64.9 73.7 35.0 40.0 29.4 40.7 60.0 75.0 40.0 56.9 50.0 40.0 48.0 45.0 47.1
Gemini 80.0 71.9 80.0 32.5 53.3 23.5 41.3 65.0 75.0 44.0 60.0 45.0 40.0 44.0 35.0 47.3
Llama3 60.0 62.2 45.0 12.5 33.3 11.8 21.1 50.0 50.0 8.0 33.9 25.0 40.0 28.0 20.0 36.6
CodeLlama 60.0 67.8 80.0 40.0 40.0 11.8 38.5 35.0 35.0 28.0 32.3 40.0 0.0 32.0 40.0 26.3
Average 77.1 75.1 79.1 39.6 44.7 40.1 47.7 64.3 70.7 40.6 57.1 52.1 42.9 50.3 51.4 51.5
l|c|cc|ccc>cccc>ccc>ccc|c|c
Original Structural Representational
Logic Alteration Concept Analysis Question Format A. Format Weighted Auto-eval
Models Seed All Gran. Reason Compute. Formul Avg. Quest. Sol. Crit. Avg. Form. Form. Avg. Form. Avg. Agreement
Samp. Samp. Adjust. Adjust. Adjust. Adjust. Perf. Under. Under. Think. Perf. Change. Comp. Perf. Constraint
MATH
o1-preview 100.0 95.2 95.0 90.0 91.7 62.9 82.2 75.0 50.0 20.0 46.2 85.0 80.0 84.0 85.0 71.9 84.3
GPT-4o 100.0 95.1 100.0 92.5 81.8 71.4 85.8 95.0 90.0 48.0 75.4 100.0 80.0 96.0 85.0 71.3 85.6
GPT-4 100.0 94.6 90.0 67.5 90.9 51.4 68.9 45.0 15.0 24.0 27.7 85.0 40.0 76.0 80.0 58.3 81.0
GPT-3.5 80.0 78.9 75.0 25.0 54.5 14.3 34.0 50.0 30.0 8.0 27.7 40.0 60.0 44.0 25.0 32.4 85.2
Gemini 80.0 91.7 80.0 60.0 81.8 31.4 56.6 45.0 10.0 12.0 21.5 45.0 40.0 44.0 70.0 45.8 86.6
Llama3 60.0 80.7 65.0 37.5 45.5 14.3 35.8 15.0 15.0 8.0 12.3 30.0 40.0 32.0 30.0 27.8 88.9
Metamath 80.0 82.3 75.0 22.5 45.5 14.3 32.1 30.0 10.0 8.0 15.4 40.0 100.0 52.0 45.0 30.6 89.8
Average 85.7 88.4 82.9 56.4 70.2 37.1 56.5 50.7 31.4 18.3 32.3 60.7 62.9 61.1 60.0 48.3 85.9
CODE
o1-preview 100.0 92.4 90.0 55.0 46.7 71.4 58.5 70.0 60.0 48.0 65.5 70.0 80.0 72.0 85.0 65.9 86.4
GPT-4o 80.0 90.2 90.0 47.5 46.7 65.7 60.9 65.0 60.0 32.0 50.7 60.0 40.0 56.0 70.0 58.2 85.5
GPT-4 80.0 76.5 100 37.5 46.7 50.0 52.3 75.0 65.0 48.0 61.5 65.0 40.0 60.0 55.0 74.5 71.7
GPT-3.5 80.0 64.9 73.7 35.0 40.0 29.4 40.7 55.0 70.0 40.0 56.9 50.0 40.0 48.0 45.0 46.3 87.2
Gemini 80.0 71.9 80.0 32.5 53.3 23.5 41.3 65.0 55.0 24.0 46.2 45.0 40.0 44.0 35.0 43.2 85.4
Llama3 60.0 62.2 45.0 12.5 33.3 11.8 21.1 50.0 40.0 0.0 46.1 25.0 40.0 28.0 25.0 36.6 91.8
CodeLlama 60.0 67.8 80.0 40.0 40.0 11.8 38.5 25.0 15.0 16.0 18.0 40.0 0.0 32.0 40.0 26.4 79.0
Average 77.1 75.1 79.8 37.1 43.8 37.7 44.8 57.9 52.1 29.7 49.3 50.7 40.0 48.6 50.7 50.2 83.9
ll|cccc|c
Models G. R. C. F. Avg.
Adj. Adj. Adj. Adj.
MATH
GPT-4 100 77.5 90.91 71.43 81.13
GPT-3.5 90 50 90.91 40 58.49
Gemini 95 57.50 63.64 45.71 61.32
CODE
GPT-4 100 50 46.67 55.88 59.43
GPT-3.5 82.35 42.50 40 26.47 43.40
Gemini 70.59 25 53.33 26.47 36.79
ll|cccc|c
Models G. R. C. F. Avg.
Adj. Adj. Adj. Adj.
Self-C
GPT-4 95 87.5 90.91 65.71 82.08
GPT-3.5 60 45 45.45 25.71 41.51
Gemini 75 45 81.82 40 52.83
POT
GPT-4 95 90 81.82 68.57 83.02
GPT-3.5 75 57.5 54.55 25.71 50
Gemini 90 60 63.64 45.71 61.32
Large Drop in Reasoning Performance of LLMs.
@@
The results show that perturbed questions significantly challenge all models in both math and coding contexts. GPT-4’s accuracy decreased notably, as did other LLMs, with all showing a performance decline of over 30 points. Notably, closed-source models outperformed open-source ones in every tested aspect. Models like CodeLlama and Metamath fine-tuned on specific tasks, performed better in logic alteration and representational perturbations but worse in concept analysis. This suggests fine-tuning may restrict broader reasoning capacities. In general, LLMs handled logic alteration better than concept analysis, indicating their robustness in abstract reasoning yet limitations in understanding deeper mathematical concepts. GPT-4 demonstrated resilience across various question types, outshining others, especially in handling different problem-solving frameworks, although it still struggled more in math than in coding in concept analysis. We include Target-wise (Level II) performance analysis in Appendix D. The latest LLMs, such as o1-preview and GPT-4o, have outperformed their predecessor, GPT-4. However, there is still a performance decline of 9% for o1-preview and 12% for GPT-4o. Notably, o1-preview demonstrates the highest resilience to perturbation tests in both Math and Coding questions, highlighting the significance of learning to reason.
5.4 Analyses
@@
In Section 5.3, providing models with the correct answer to the original question along with the prompt significantly improves their ability to solve perturbed questions, particularly in the Computational Adjustment dimension. However, performance remains weak in Formula Adjustment, highlighting challenges in abstract reasoning despite access to solutions. Notably, even equipped with correct answers, some models like Gemini and GPT-3.5 still fail on simpler question variants, underscoring their low sensitivity to semantic perturbations.
@@
In Section 5.3, different prompting techniques greatly influence model performance in Logic Alteration tasks. The Program-of-Thoughts technique notably boosts reasoning capabilities in closed-source models by reducing logical errors, leading to better performance in symbolic manipulation for GPT-4. Conversely, the Self-Consistency method shows only minor improvements and even a performance decline in the Gemini model, suggesting difficulties in effective in-context learning for new unseen tasks.
@@
The Formulation Adjustment dimension presents a significant challenge to both closed-source and open-source models, largely due to the demands of abstract reasoning. Instead of reasoning a number or code as the answer, this involves the manipulation of abstract math and coding concepts in the logical space behind the surface of the problem. For example, in the WhatIf category, models must hypothesize outcomes by changing certain events under consistent conditions, which requires a nuanced grasp of the problem-solving framework. The Critical Thinking dimension tests a model’s ability to scrutinize relationships between pieces of information, demanding a comprehensive analysis to identify inconsistencies without a predefined solution path. This emphasizes the necessity for models to thoroughly understand and navigate through all possible avenues to effectively resolve conflicts or discrepancies. Furthermore, the Format Change dimension poses difficulties to models like GPT-3.5 attempt to follow these constraints but often fail to maintain the integrity of their reasoning when adapting to new formats, highlighting a lack of flexibility in handling varied task demands.
@@ 6.1 Difficulty Change by Perturbations
l|ccc
Category Human Acc() Model Acc() Time Consumption
Number of Reasoning Steps
95.2(-4.8) 32.0(-44.0) 178%
98.4(-1.6) 44.4(-31.6) 39%
Reasoning Depth
97.3(-2.7) 38.4(-37.6) 113%
97.7(-2.3) 42.1(-33.9) 62%
@@ 7 Conclusion
Our study evaluated the robustness of several prominent Large Language Models (LLMs) in handling mathematical and coding problems. By employing an ontology for random perturbations on questions from the GSM8K and HumanEval datasets, we crafted two specialized datasets, GSMore and HumanEval-Core, containing 216 and 219 questions respectively. These datasets target a broad variations of mathematical and coding problem-solving and analytical skills, resulting in notable performance drops in LLMs upon evaluation. The introduction of GSMore and HumanEval-Core provides a new framework for assessing LLMs’ abilities in mathematics and coding, while also revealing their vulnerabilities in consistent reasoning across different formats. This research highlights the complex challenges that LLMs face, stressing the importance of continued exploration into their strengths and weaknesses in logical reasoning tasks. Our dataset GSMore 222https://huggingface.co/datasets/declare-lab/GSM8k_MORE and HumanEval-Core 333https://huggingface.co/datasets/declare-lab/HumanEval_CORE will be publicly available online.